OpenStudy (anonymous):

If x=2 Solve for x: x+8 @shrutipande9

OpenStudy (cp9454):

10

OpenStudy (anonymous):

o_O ten?

OpenStudy (anonymous):

Oh wait, sorry. x+8=5

OpenStudy (cp9454):

x= -3

OpenStudy (anonymous):

How you do that.

OpenStudy (cp9454):

x+8 = 5 x = 5-8 x = -3

OpenStudy (tester97):

This would work better if you used latex then explained your steps @cp9454

OpenStudy (cp9454):

i didn't knew that, how to use that?

OpenStudy (tester97):

http://prntscr.com/40zvni Click this button and write out your math equations here http://prntscr.com/40zw6u

OpenStudy (cp9454):

thank u @tester97

OpenStudy (anonymous):

I just don't how to do that. I've tried everything. I assumed x as ket vector <x| then found its eigenvalue by using a hermitian matrix to find lambda then applied Schrödinger's time-dependent variable to find the given state of the qubit: \[i \hbar \frac{ \partial }{ \partial t }\psi (r,t) = [\frac{ -\hbar }{ 2m} gradient^2 +V(r,t)]\Psi (r,t)\]

OpenStudy (tester97):

welcome :)

OpenStudy (anonymous):

Then I applied U-substitution to find the initial state of a qubit before it is observed in a plus/minus basis and then used Grover's algorithm to search through a lattice of N-Indexes.

OpenStudy (anonymous):

I'm not following here...

OpenStudy (anonymous):

Just hold on... let's take this back a few steps... How do I start this?

OpenStudy (anonymous):

Jesus Christ... can anyone help me with this? I'm trying to finish my homework :(

OpenStudy (cp9454):

first explain your question or write it properly.

OpenStudy (anonymous):

If x=2 Solve for x: x+8=5

OpenStudy (cp9454):

look it is given that x=2, and from given equation we are getting x=-3 since, \[-3 \neq 2 \] therefore \[x \in \phi\]. means no solution.

OpenStudy (anonymous):

How you do that

OpenStudy (tester97):

Dont forget to medal people for all their hard work :)

OpenStudy (anonymous):

I'm not following :( I gather that first you would find it's Eigenvalue by using a transposed hermitian matrix: \[H(dis.)+|x>=\lambda + |x> \] Correct?